OpenStudy (anonymous):

(3x10^-5)(3x10^-1) write it in standard notation

OpenStudy (anonymous):

multiply \(3\times 3\) and add the exponents

OpenStudy (anonymous):

\[(3\times 10^{-5})(3\times 10^{-1})=3\times 3\times 10^{-5+(-1)}=9\times 10^{-6}\]

OpenStudy (anonymous):

now if you want "standard notation, move the decimal point to the left six places